What is the mole fraction of the KCl solute in a 1.00 molal aqueous solution?
1
Expert's answer
2020-12-18T08:47:08-0500

"n = \frac{m}{M}"

"M(H_2O) = 18\ \frac{g}{mol}"

"m = 1\ kg; n(H_2O) = \frac{1000\ g}{18\ g/mol} \approx 55.56\ mol"

1 molal means 1 mol of solute in 1 kg of a solvent

"X_{KCl} = \frac{1\ mol}{1\ mol + 55.56\ mol} =\frac{1}{56.56} \approx 0.01768"


Answer: "X_{KCl} = 0.01768"